Given a Vmax of 7.45V and a modulation index of 0.691, calculate for Vmin

1
Expert's answer
2021-10-01T00:03:44-0400
"m=\dfrac{V_{max}-V_{min}}{V_{max}+V_{min}}"

"m(V_{max}+V_{min})=V_{max}-V_{min}"

"V_{min}=\dfrac{1-m}{1+m}V_{max}"

"V_{min}=\dfrac{1-0.691}{1+0.691}(7.45\ V)=1.36\ V"

"V_{min}=1.36\ V"